# The role of turbulence in setting the phase of the ISM and implications for the star formation rate

### v2

@@ 1 Introduction
The interstellar medium (ISM) consists mainly of hydrogen gas which occurs in ionised (HII), atomic (HI) and molecular (H2) form. It has been established that HI itself has two thermal phases that coexist at the same pressure: the warm neutral medium (WNM) with temperatures of a few thousand Kelvin and the cold neutral medium (CNM) with temperatures below 250 K (Field et al., 1969; Wolfire et al., 1995, 2003; McClure-Griffiths et al., 2023). Gas in one of these two regimes is thermally stable, meaning that a small perturbation cannot change its temperature dramatically. On the other hand, gas with a temperature in the range between the CNM and WNM temperatures is thermally unstable and could undergo rapid transition to either the WNM or CNM state. This regime is called the unstable or lukewarm neutral medium (UNM or LNM).
@@
Gas is the main raw material from which stars form. As such, relations exist between the gas reservoir and the star formation rate (SFR), which can be averaged over full galaxies or sub-regions within, or measured for individual star-forming clouds. Schmidt (1959) proposed a power-law relation between the gas density and the SFR. The relation between the surface density, a quantity which is more directly observable especially for extra-galactic measurements, and SFR is commonly referred to as the Kennicutt-Schmidt (KS) relation (Kennicutt, 1998; Kennicutt & Evans, 2012). On average, it has the form SFR $\propto\Sigma_{\mathrm{HI+H2}}^{1.4}$ above 9 $\mathrm{M}_{\odot}$pc-2, where the relation exhibits a break, and below which it shows steeper slopes and more scatter (Kennicutt, 1998; Bigiel et al., 2008).
@@
The relation becomes even more direct when considering only the molecular gas. Bigiel et al. (2008) have established a tight linear relation $\Sigma_{\mathrm{H}_{2}}\propto\Sigma_{\mathrm{SFR}}$, where H2 was estimated from CO emission (see also Pessa et al., 2021; Sun et al., 2023). In contrast, the relation between the SFR and the HI surface density is found to vary from galaxy to galaxy (Bigiel et al., 2008; Leroy et al., 2008; Shetty et al., 2014). Since only cold gas is relevant for the formation of stars (Krumholz et al., 2009a; Klessen & Glover, 2016), regardless of whether it is in atomic or molecular form (Glover & Clark, 2012), we would not expect a universal relation of the SFR with the total amount of HI if the CNM fraction varies with environment.
@@
Decomposition of HI into WNM and CNM is a difficult task and so far observational results are limited to the Milky Way or other Local Group galaxies (Dickey, 1979; Dickey et al., 2000; HI4PI Collaboration et al., 2016; Murray et al., 2018). In their recent review, McClure-Griffiths et al. (2023) compiled an overview of CNM fraction measurements in the Galaxy. A large variety is recovered, with values ranging from zero up to 80% or more. Several trends with environment appear. Heiles & Troland (2003) found a CNM fraction that on average increases with total HI column density up to $\Sigma_{\mathrm{HI}}\approx 9$ $\mathrm{M}_{\odot}$pc-2, which marks the threshold between HI and H2 dominated gas at solar metallicity (Bigiel et al., 2008; Leroy et al., 2008). Observations near molecular clouds typically show relatively high CNM fractions, highlighting the important connection between these two gas phases. Gas at high Galactic latitudes is found to have lower CNM fractions (Kalberla & Kerp, 2009; Marchal & Miville-Deschênes, 2021).
@@
The star formation process in galaxies can thus be thought of as a sequence of steps: the condensation of CNM out of WNM followed by the possible formation of a molecular cloud out of this CNM, in which then finally dense collapsing cores form and produce stars (Schruba et al., 2011; Smith et al., 2023). Star formation can potentially be regulated through either one of these processes and the dominant limiting process is likely dependent on environment.
@@
We know that turbulence plays a key role in star formation (e.g. Mac Low & Klessen, 2004; McKee & Ostriker, 2007). Observations show that molecular clouds are turbulent and its importance is now widely accepted (Roman-Duval et al., 2010; Hennebelle & Falgarone, 2012; Miville-Deschênes et al., 2017). Analytical theories and simulations show that turbulence interacts with gravity in two ways. On the one hand, it widens the density PDF, creating more overdensities that may become gravitationally unstable, which is favourable for star formation. On the other hand, it provides support against gravitational collapse in the form of turbulent velocity dispersion, which reduces the rate at which dense gas is transformed into stars. These principles are the basis of the theory of turbulent fragmentation (Krumholz & McKee, 2005; Padoan & Nordlund, 2011; Hennebelle & Chabrier, 2011; Federrath & Klessen, 2012; Hennebelle et al., 2024).
@@
It is thought that the turbulence in molecular clouds is inherited from larger scales during their formation process (Heyer & Brunt, 2004; Brunt et al., 2009; Klessen & Hennebelle, 2010). What exactly drives the turbulence on these scales is still debated. In present-day spiral galaxies such as the Milky Way, numerical studies indicate that stellar feedback in the form of supernova (SN) possibly in combination with ionising radiation and winds is the dominant driving mechanism in most parts of the galaxy (Peters et al., 2017; Gatto et al., 2017; Colling et al., 2018; Rathjen et al., 2021). These studies reproduce levels of star formation in agreement with observations, without needing to invoke additional limiting processes. Recently, Colman et al. (2024) investigated simulated cloud properties in different environments and from different numerical setups. They found an overall trend of the turbulent velocity dispersion consistent with large-scale turbulent decay. Furthermore, at small spatial scales – and small molecular clouds – nearby SNe leave a noticeable imprint on the velocity dispersion.
@@
The situation may be different in the centre and outskirts of galaxies. Sun et al. (2020) showed that the velocity dispersion of clouds in the centre of barred spiral galaxies is elevated. The same is true for the central region of our Milky Way (Shetty et al., 2012). The fact that these regions lie below the KS relation (Longmore et al., 2013), leads us to speculate that injection of turbulence from galactic dynamics plays a role here. Also in galaxies which have high gas densities, such as found at high redshift, large-scale galactic processes may be important drivers of turbulence (Krumholz et al., 2018). In their kiloparsec box simulations, Brucy et al. (2020, 2023) found that large-scale driving quenches star formation for column densities above 20 $\mathrm{M}_{\odot}$pc-2 and that the KS relation cannot be reproduced with stellar feedback alone.
@@ 2 Simulations
As in previous work (Iffrig & Hennebelle, 2015; Colling et al., 2018; Brucy et al., 2020), we model a 1 kpc3 region of a galaxy with realistic ISM physics. The first set of simulations used for this work are described in Colman et al. (2022). Additionally, we extend this set to include four new runs with lower initial column density, to probe a wider variety of conditions. What is characteristic about these simulations is that they assume a constant UV background. To gain additional insight, we analysed a second set of simulations with a variable UV field. The details of the first set are described below. In Section 5.1, we discuss the differences with the second set.
@@ 2.2 Numerics
To evolve our simulation in time, we used the adaptive mesh refinement magneto-hydrodynamics code ramses (Teyssier, 2002; Fromang et al., 2006), with treatments of the magnetic field in the ideal approximation and radiation using a moment-based method (Rosdahl et al., 2013). The simulation box is periodic in the $x$- and $y$-direction and has open boundary conditions in the $z$-direction. The coarse grid has a resolution of 3.9 pc and is further refined based on a mass criterion up to a maximum resolution of 0.24 pc in the densest regions. Aside from self-gravity, we also apply an external gravitational potential representing the effects of the stellar disk as prescribed by Kuijken & Gilmore (1989); Joung & Mac Low (2006). We use the ISM cooling/heating model from Audit & Hennebelle (2005) which is based on the work of Wolfire et al. (2003), further detailed in one of the next subsections. This model produces a multi-phase ISM with a warm and cold phase in our simulations, as demonstrated later on in this paper.
@@ 2.3 Sink particles and stellar feedback
When the gravitational collapse reaches the resolution limits, we introduce sink particles (Bate et al., 1995; Federrath et al., 2010b) according to the recipe prescribed by Bleuler & Teyssier (2014). After their birth, sinks accrete gas according to the threshold accretion scheme: gas which is above the sink formation threshold of $10^{4}$ cm-3 and within the accretion radius of 4 cells will be accreted. The high-$\Sigma$ simulations have been evolved for 60 Myr during which 2 to 5 $\%$ of the gas has been converted into stars. The low-$\Sigma$ runs are computationally less demanding and have varying final times, depending on the intensity of their star formation activity.
Each time a sink has accreted 120 $\mathrm{M}_{\odot}$of gas, it forms an individual massive star with a mass between 8 and 120 $\mathrm{M}_{\odot}$, randomly drawn from the Salpeter IMF (Salpeter, 1955). The remainder of the mass in the sink particle is considered to be distributed between low mass stars. The total mass of a sink is thus equal to the sum of all individual massive stars that formed in this sink, in addition to the mass of an unresolved collection of low-mass stars.
The massive stars are sources of stellar feedback. They emit ionising radiation from the position of the sink particle in which they formed and explode as a supernova at the end of their lifetime at a random location around the sink (Rosdahl et al., 2013; Geen et al., 2016; Iffrig & Hennebelle, 2017; Colling et al., 2018). The supernovae inject $4\times 10^{43}$ g cm s-1 momentum and $10^{51}$ erg of thermal energy, but the velocity and temperature of the surrounding gas are limited to 300 km s-1 and $10^{6}$ K respectively to prevent the time step from becoming very small. We have tested the impact of this procedure by comparing several statistics of runs with and without these restrictions and find limited differences. A comparison regarding the SFR has been presented in Brucy et al. (2023). In the Appendix B, we additionally show the velocity dispersion as a function of density. We find that the limitations do not significantly affect the results. The physical reason is that when very high velocities (typically larger than 300 km s-1) are needed, it is because the supernova explodes in a very tenuous medium. Such explosions have very limited impact on the dense gas, as studied for instance by Iffrig & Hennebelle (2015). For low-mass stars, the UV output is negligible and their lifetimes are much longer than the simulation time. Due to their computational cost, we do not include stellar winds. Furthermore, similar numerical studies found that radiation is more important for the regulation of the SFR and that including winds further reduces the SFR by only a factor 2 (Rathjen et al., 2021).
@@ 2.4 Turbulence driving
In this work, we study the effect of turbulent energy injection from large scales. This presents a source of turbulence additional to stellar feedback, mimicking the possible effect of galactic dynamical processes that occur on scales larger than the box size, such as spiral arms, a bar, or interactions with neighbouring galaxies. A detailed prescription of the turbulence driving can be found in Colman et al. (2022) or Brucy et al. (2023). In short, a bi-dimensional turbulence driving force aligned with the disk plane is added as an additional stochastic external force in the Euler equation. This driving acts on scales between a full and a third of the box length and is mainly solenoidal, with a solenoidal fraction of 0.75. We explore several different forcing strengths, ranging from very weak to strong, as well as running reference simulations without external forcing (see Table 1).
In Figure 1 we compare the energy injected by large-scale driving to an estimate of the turbulence injection by SNe per unit time for each simulation. We count an energy of 1051 erg at an average efficiency of 6% per SN. Note that this efficiency is difficult to quantify and it depends on the environment in which the SN goes off (Martizzi et al., 2015; Iffrig & Hennebelle, 2017). After an initialisation phase during which the SN rate increases, the feedback self-regulated to a fairly stable rate which is mostly dependent on the column density of the simulation. Figure 1 serves as a guide for interpreting the relevance of both driving mechanisms in the remainder of this paper.
Different forcing strengths lead to different final global velocity dispersions. In Figure 2, we show the evolution of the velocity dispersion in the different simulations, measured in the disk. This quantity reflects the differences in energy injection rate from Figure 1. We note that the values obtained in some of the low-$\Sigma$ runs are only around 2-3 km s-1, which is significantly lower than what is usually inferred in simulations and observations. This is due to the low SN rate in these runs. We will come back to this in Section 5.4. For the other simulations, our measured velocity dispersions are very similar to what is reported in other work, for instance Girichidis et al. (2016) and Kim et al. (2023) (their Fig. 19 and Table 5 respectively).
@@ 2.5 Cooling and heating model
What sets the thermal balance in the ISM is a topic that has been well-studied in the past decades. Wolfire et al. (1995) summarised which processes are relevant and provide a model for the solar neighbourhood, which was later extended for other Galactic radii (Wolfire et al., 2003). More recent improvements include a more accurate treatment for low-metallicity environments (Glover & Clark, 2014; Bialy & Sternberg, 2019). In the context of this work, we use the ISM cooling/heating model from Audit & Hennebelle (2005). We assume solar metallicity and solar neighborhood abundances. We briefly review the dominant processes included in this model.
The dominant heating process is photoelectric heating from small dust grains and PAHs, that is the emission of electrons from the grains after exposure to the interstellar radiation field. While UV radiation can free electrons in grains of all sizes, these electrons can escape more easily from small grains. Consequently, the photoelectric heating rate depends on the amount of small grains available and the strength of the far UV field $G_{0}$. Bakes & Tielens (1994) provide an analytical model as a function of $G_{0}$, the temperature $T$ and the electron density $n_{e}$, assuming a fixed grain size distribution. Even though the photoionisation of hydrogen is followed using Ramses-RT, the electron density used in the cooling model is calculated using the formula provided by Wolfire et al. (2003). Corrections for photoionisation in the WNM are made based on the work of Ferland (2003), as is done in Geen et al. (2016). A constant uniform UV background with a strength equal to 0.6 Habing units is assumed (Habing, 1969). In reality, we expect the UV field to vary locally and to be proportional to the local SFR (Ostriker et al., 2010). The advantage of keeping the UV background constant and uniform is that it makes it easy to disentangle between the dynamical effects, induced by the stellar feedback and turbulent driving, and the thermal effect induced by UV heating. In Section 5, we will explore a second set of simulations where the UV heating is time-dependent and proportional to the star formation rate (but still uniform spatially).

### v3

@@ 1 Introduction
The interstellar medium (ISM) consists mainly of hydrogen gas which occurs in ionised (HII), atomic (HI) and molecular (H2) form. It has been established that HI itself has two thermal phases that coexist at the same pressure: the warm neutral medium (WNM) with temperatures of a few thousand Kelvin and the cold neutral medium (CNM) with temperatures below 250 K (Field_et_al1969; Wolfire_et_al1995; Wolfire_et_al2003; McClure-Griffiths_et_al2023). Gas in one of these two regimes is thermally stable, meaning that a small perturbation cannot change its temperature dramatically. On the other hand, gas with a temperature in the range between the CNM and WNM temperatures is thermally unstable and could undergo rapid transition to either the WNM or CNM state. This regime is called the unstable or lukewarm neutral medium (UNM or LNM).
@@
Gas is the main raw material from which stars form. As such, relations exist between the gas reservoir and the star formation rate (SFR), which can be averaged over full galaxies or sub-regions within, or measured for individual star-forming clouds. Schmidt1959 proposed a power-law relation between the gas density and the SFR. The relation between the surface density, a quantity which is more directly observable especially for extra-galactic measurements, and SFR is commonly referred to as the Kennicutt-Schmidt (KS) relation (Kennicutt1998; Kennicutt&Evans2012). On average, it has the form SFR $\propto\Sigma_{\mathrm{HI+H2}}^{1.4}$ above 9 $\mathrm{M}_{\odot}$pc-2, where the relation exhibits a break, and below which it shows steeper slopes and more scatter (Kennicutt1998; Bigiel_et_al2008).
@@
The relation becomes even more direct when considering only the molecular gas. Bigiel_et_al2008 have established a tight linear relation $\Sigma_{\mathrm{H}_{2}}\propto\Sigma_{\mathrm{SFR}}$, where H2 was estimated from CO emission (see also Pessa2021; Sun2023). In contrast, the relation between the SFR and the HI surface density is found to vary from galaxy to galaxy (Bigiel_et_al2008; Leroy_et_al2008; Shetty2014). Since only cold gas is relevant for the formation of stars (Krumholz_et_al2009; Klessen2016), regardless of whether it is in atomic or molecular form (Glover2012), we would not expect a universal relation of the SFR with the total amount of HI if the CNM fraction varies with environment.
@@
Decomposition of HI into WNM and CNM is a difficult task and so far observational results are limited to the Milky Way or other Local Group galaxies (Dickey1979; Dickey_et_al2000; HI4PI2016; Murray_et_al2018). In their recent review, McClure-Griffiths_et_al2023 compiled an overview of CNM fraction measurements in the Galaxy. A large variety is recovered, with values ranging from zero up to 80% or more. Several trends with environment appear. Heiles_Troland2003 found a CNM fraction that on average increases with total HI column density up to $\Sigma_{\mathrm{HI}}\approx 9$ $\mathrm{M}_{\odot}$pc-2, which marks the threshold between HI and H2 dominated gas at solar metallicity (Bigiel_et_al2008; Leroy_et_al2008). Observations near molecular clouds typically show relatively high CNM fractions, highlighting the important connection between these two gas phases. Gas at high Galactic latitudes is found to have lower CNM fractions (Kalberla_Kerp2009; Marchal_Miville-Deschenes2021).
@@
The star formation process in galaxies can thus be thought of as a sequence of steps: the condensation of CNM out of WNM followed by the possible formation of a molecular cloud out of this CNM, in which then finally dense collapsing cores form and produce stars (Schruba_et_al2011; Smith_at_al2023). Star formation can potentially be regulated through either one of these processes and the dominant limiting process is likely dependent on environment.
@@
We know that turbulence plays a key role in star formation (e.g. MacLow2004; McKee2007). Observations show that molecular clouds are turbulent and its importance is now widely accepted (Roman-Duval_et_al2010; Hennebelle_Falgarone2012; Mivilledeschenes_et_al2017). Analytical theories and simulations show that turbulence interacts with gravity in two ways. On the one hand, it widens the density PDF, creating more overdensities that may become gravitationally unstable, which is favourable for star formation. On the other hand, it provides support against gravitational collapse in the form of turbulent velocity dispersion, which reduces the rate at which dense gas is transformed into stars. These principles are the basis of the theory of turbulent fragmentation (Krumholz_McKee2005; Padoan_Nordlund2011; Hennebelle_Chabrier2011; Federrath2012; Hennebelle_et_al2024).
@@
It is thought that the turbulence in molecular clouds is inherited from larger scales during their formation process (Heyer2004; Brunt_et_al2009; Klessen2010). What exactly drives the turbulence on these scales is still debated. In present-day spiral galaxies such as the Milky Way, numerical studies indicate that stellar feedback in the form of supernova (SN) possibly in combination with ionising radiation and winds is the dominant driving mechanism in most parts of the galaxy (Peters_et_al2017; Gatto_et_all2017; Colling_et_al2018; Rathjen_et_al2021). These studies reproduce levels of star formation in agreement with observations, without needing to invoke additional limiting processes. Recently, colman24 investigated simulated cloud properties in different environments and from different numerical setups. They found an overall trend of the turbulent velocity dispersion consistent with large-scale turbulent decay. Furthermore, at small spatial scales – and small molecular clouds – nearby SNe leave a noticeable imprint on the velocity dispersion.
@@
The situation may be different in the centre and outskirts of galaxies. Sun_et_al2020 showed that the velocity dispersion of clouds in the centre of barred spiral galaxies is elevated. The same is true for the central region of our Milky Way (Shetty2012). The fact that these regions lie below the KS relation (Longmore2013), leads us to speculate that injection of turbulence from galactic dynamics plays a role here. Also in galaxies which have high gas densities, such as found at high redshift, large-scale galactic processes may be important drivers of turbulence (Krumholz_et_al2018). In their kiloparsec box simulations, Brucy_et_al2020; Brucy_et_al2023 found that large-scale driving quenches star formation for column densities above 20 $\mathrm{M}_{\odot}$pc-2 and that the KS relation cannot be reproduced with stellar feedback alone.
@@ 2 Simulations
As in previous work (Iffrig&Hennebelle2015; Colling_et_al2018; Brucy_et_al2020), we model a 1 kpc3 region of a galaxy with realistic ISM physics. The first set of simulations used for this work are described in colman22. Additionally, we extend this set to include four new runs with lower initial column density, to probe a wider variety of conditions. What is characteristic about these simulations is that they assume a constant UV background. To gain additional insight, we analysed a second set of simulations with a variable UV field. The details of the first set are described below. In Section 5.1, we discuss the differences with the second set.
@@ 2.2 Numerics
To evolve our simulation in time, we used the adaptive mesh refinement magneto-hydrodynamics code ramses (Teyssier2002; Fromang_et_al2006), with treatments of the magnetic field in the ideal approximation and radiation using a moment-based method (Rosdahl_et_al2013). The simulation box is periodic in the $x$- and $y$-direction and has open boundary conditions in the $z$-direction. The coarse grid has a resolution of 3.9 pc and is further refined based on a mass criterion up to a maximum resolution of 0.24 pc in the densest regions. Aside from self-gravity, we also apply an external gravitational potential representing the effects of the stellar disk as prescribed by Kuijken_Gilmore1989; Joung_MacLow2006. We use the ISM cooling/heating model from Audit_Hennebelle2005 which is based on the work of Wolfire_et_al2003, further detailed in one of the next subsections. This model produces a multi-phase ISM with a warm and cold phase in our simulations, as demonstrated later on in this paper.
@@ 2.3 Sink particles and stellar feedback
When the gravitational collapse reaches the resolution limits, we introduce sink particles (Bate_et_al1995; Federrath_et_al2010) according to the recipe prescribed by Bleuler&Teyssier2014. After their birth, sinks accrete gas according to the threshold accretion scheme: gas which is above the sink formation threshold of $10^{4}$ cm-3 and within the accretion radius of 4 cells will be accreted. The high-$\Sigma$ simulations have been evolved for 60 Myr during which 2 to 5 $\%$ of the gas has been converted into stars. The low-$\Sigma$ runs are computationally less demanding and have varying final times, depending on the intensity of their star formation activity.
Each time a sink has accreted 120 $\mathrm{M}_{\odot}$of gas, it forms an individual massive star with a mass between 8 and 120 $\mathrm{M}_{\odot}$, randomly drawn from the Salpeter IMF (Salpeter1955). The remainder of the mass in the sink particle is considered to be distributed between low mass stars. The total mass of a sink is thus equal to the sum of all individual massive stars that formed in this sink, in addition to the mass of an unresolved collection of low-mass stars.
The massive stars are sources of stellar feedback. They emit ionising radiation from the position of the sink particle in which they formed and explode as a supernova at the end of their lifetime at a random location around the sink (Rosdahl_et_al2013; Geen_et_al2016; Iffrig&Hennebelle2017; Colling_et_al2018). The supernovae inject $4\times 10^{43}$ g cm s-1 momentum and $10^{51}$ erg of thermal energy, but the velocity and temperature of the surrounding gas are limited to 300 km s-1 and $10^{6}$ K respectively to prevent the time step from becoming very small. We have tested the impact of this procedure by comparing several statistics of runs with and without these restrictions and find limited differences. A comparison regarding the SFR has been presented in Brucy_et_al2023. In the Appendix B, we additionally show the velocity dispersion as a function of density. We find that the limitations do not significantly affect the results. The physical reason is that when very high velocities (typically larger than 300 km s-1) are needed, it is because the supernova explodes in a very tenuous medium. Such explosions have very limited impact on the dense gas, as studied for instance by Iffrig&Hennebelle2015. For low-mass stars, the UV output is negligible and their lifetimes are much longer than the simulation time. Due to their computational cost, we do not include stellar winds. Furthermore, similar numerical studies found that radiation is more important for the regulation of the SFR and that including winds further reduces the SFR by only a factor 2 (Rathjen_et_al2021).
@@ 2.4 Turbulence driving
In this work, we study the effect of turbulent energy injection from large scales. This presents a source of turbulence additional to stellar feedback, mimicking the possible effect of galactic dynamical processes that occur on scales larger than the box size, such as spiral arms, a bar, or interactions with neighbouring galaxies. A detailed prescription of the turbulence driving can be found in colman22 or Brucy_et_al2023. In short, a bi-dimensional turbulence driving force aligned with the disk plane is added as an additional stochastic external force in the Euler equation. This driving acts on scales between a full and a third of the box length and is mainly solenoidal, with a solenoidal fraction of 0.75. We explore several different forcing strengths, ranging from very weak to strong, as well as running reference simulations without external forcing (see Table 1).
In Figure 1 we compare the energy injected by large-scale driving to an estimate of the turbulence injection by SNe per unit time for each simulation. We count an energy of 1051 erg at an average efficiency of 6% per SN. Note that this efficiency is difficult to quantify and it depends on the environment in which the SN goes off (Martizzi_et_al2015; Iffrig&Hennebelle2017). After an initialisation phase during which the SN rate increases, the feedback self-regulated to a fairly stable rate which is mostly dependent on the column density of the simulation. Figure 1 serves as a guide for interpreting the relevance of both driving mechanisms in the remainder of this paper.
Different forcing strengths lead to different final global velocity dispersions. In Figure 2, we show the evolution of the velocity dispersion in the different simulations, measured in the disk. This quantity reflects the differences in energy injection rate from Figure 1. We note that the values obtained in some of the low-$\Sigma$ runs are only around 2-3 km s-1, which is significantly lower than what is usually inferred in simulations and observations. This is due to the low SN rate in these runs. We will come back to this in Section 5.4. For the other simulations, our measured velocity dispersions are very similar to what is reported in other work, for instance girichidis2016 and kimchang2023 (their Fig. 19 and Table 5 respectively).
@@ 2.5 Cooling and heating model
What sets the thermal balance in the ISM is a topic that has been well-studied in the past decades. Wolfire_et_al1995 summarised which processes are relevant and provide a model for the solar neighbourhood, which was later extended for other Galactic radii (Wolfire_et_al2003). More recent improvements include a more accurate treatment for low-metallicity environments (Glover_Clark2014; Bialy_Sternberg2019). In the context of this work, we use the ISM cooling/heating model from Audit_Hennebelle2005. We assume solar metallicity and solar neighborhood abundances. We briefly review the dominant processes included in this model.
The dominant heating process is photoelectric heating from small dust grains and PAHs, that is the emission of electrons from the grains after exposure to the interstellar radiation field. While UV radiation can free electrons in grains of all sizes, these electrons can escape more easily from small grains. Consequently, the photoelectric heating rate depends on the amount of small grains available and the strength of the far UV field $G_{0}$. Bakes_Tielens1994 provide an analytical model as a function of $G_{0}$, the temperature $T$ and the electron density $n_{e}$, assuming a fixed grain size distribution. Even though the photoionisation of hydrogen is followed using Ramses-RT, the electron density used in the cooling model is calculated using the formula provided by Wolfire_et_al2003. Corrections for photoionisation in the WNM are made based on the work of Ferland2003, as is done in Geen_et_al2016. A constant uniform UV background with a strength equal to 0.6 Habing units is assumed (Habing1969). In reality, we expect the UV field to vary locally and to be proportional to the local SFR (Ostriker_et_al2010). The advantage of keeping the UV background constant and uniform is that it makes it easy to disentangle between the dynamical effects, induced by the stellar feedback and turbulent driving, and the thermal effect induced by UV heating. In Section 5, we will explore a second set of simulations where the UV heating is time-dependent and proportional to the star formation rate (but still uniform spatially).